OpenStudy (anonymous):

initial value problem: \[y y'=x\textrm{e}^{-y^2}\qquad y(0)=2\]

OpenStudy (anonymous):

\[y y'=x\textrm{e}^{-y^2}\qquad y(0)=2\]

OpenStudy (anonymous):

What type of equation is this?

OpenStudy (anonymous):

hell if I know, short of a differential equation.

OpenStudy (anonymous):

You want to do whats called separation of variables, im writing it out on paper, i'll post in a sec.

OpenStudy (anonymous):

so far ive got \[e^{y^2}=x^2\]

OpenStudy (anonymous):

\[y=\sqrt(2\ln(x))\]

OpenStudy (anonymous):

Dang it, I always complicate things! I was thinking it was some kind of Bernoulli lol

OpenStudy (anonymous):

Forgot constants. That is quite impressive you scan a handwritten working.

OpenStudy (anonymous):

Sometimes its faster for me than typing in LaTeX lol.

OpenStudy (anonymous):

I got \[C = e^4\] and \[y = sqrt(2ln(x^2+e^4))\]

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

i dont think that 2 belongs there.

OpenStudy (vijay):

\[\sqrt{\log(x^{2}} + e ^{4})\]

OpenStudy (anonymous):

lol you are correct. I pulled it down from the log but then . . . didnt. w/e answer is correct now. Thanks!